# Optimistic Task Inference for Behavior Foundation Models

### v1

@@ A.5.1 Determinants
A.5.2 Elliptical Potential Lemma
@@ A.5.3 UCB Optimisation
$=\hat{z}_{t}^{\top}x+\max_{\left\lVert{u}\right\rVert_{V_{t-1}}\leq\beta_{t}}u^{\top}\psi(s_{t},z)$ (84)
$\leq\hat{z}_{t}^{\top}\psi(s_{t},z)+\max_{\left\lVert{u}\right\rVert_{V_{t-1}}\leq\beta_{t}}\left\lVert{u}\right\rVert_{V_{t-1}}\left\lVert{\psi(s_{t},z)}\right\rVert_{V_{t-1}^{-1}}$ (85)
$=\hat{z}_{t}^{\top}\psi(s_{t},z)+\beta_{t}\left\lVert{\psi(s_{t},z)}\right\rVert_{V_{t-1}^{-1}}$ (86)
@@ B.2 Requesting Reward Labels Explicitly
This approach can be directly applied to OpTI-BFM by introducing a hyper-parameter $\kappa$ and only requesting $r_{t}$ if $\Delta(\phi_{t})\geq\kappa$. We evaluate labeling efficiency for this variant: for each labeling budget $n$ (# Samples), we measure average rewards at the time step $t$ in which the budget is exhausted, i.e. $\tfrac{1}{t}\sum_{i=1}^{t}r_{t}$. This is presented in Fig.˜7 (see also Fig.˜12 and Fig.˜13 for per-task results). We observe that $\kappa$ trades-off environment interaction and labeling cost: for higher $\kappa$, good performance is achieved with fewer labels, but more episodes are required to request the same number of labels. Given a specific cost for labels, a slightly more involved thresholding scheme (Tucker et al., 2023) could potentially be used to decide the threshold apriori. Surprisingly, this variant of OpTI-BFM can maintain the performance of the main algorithm, especially in easier tasks, while reducing the amount of labeled data by one order of magnitude.
@@
Figure 7: Average reward in Cheetah at different numbers of requested reward labels (# Samples) of our main methods for different information thresholds $\kappa$. We stop interaction after 30k environment steps. $\kappa$ trades-off interaction cost with labeling cost. More than one order of magnitude less reward labels can still result in the same performance in easier tasks! We report per method results of all environments in Fig.˜13 and Fig.˜12
@@ Appendix E Experiment Protocol
We evaluate our methods on a common zero-shot RL benchmark (Touati et al., 2023; Agarwal et al., 2025b). The benchmark consists of the Walker, Cheetah, and Quadruped environments from DMC (Tassa et al., 2018) with four tasks (reward functions) each. Note that each task has randomized initial states, so when evaluating single episode performance, e.g. Fig.˜9, we report the mean over 20 episodes. And when evaluating task inference performance over multiple episode, e.g., Figs.˜2 and 10, we report the mean over 10 trials. All error-bars and shaded regions denote min-max-intervals over three training seeds around mean performance.
@@ F.1 OpTI-BFM
We found both OpTI-BFM and OpTI-BFM-TS to be very robust to the hyper-parameters we tested. For each method we consider a single hyper-parameter with three values each: For UCB, we test a fixed $\beta_{t}=\beta\in\{1.0,0.1,0.001\}$. For TS, we test $\sigma\in\{0.1,0.001,0.0001\}$. This means all other parameters where held constant. Specifically, $\rho=1$ and $\lambda=1$ if not reported otherwise (Fig.˜5). The number of samples for the UCB optimization was $n=128$ throughout all experiments.
@@ F.2 LoLA
  • horizon length and task embedding update rate $\{50,100,250\}$ as in Sikchi et al. (2025);
  • learning rate $\{0.1,0.05\}$ as in Sikchi et al. (2025);

### v2

@@ A.6.1 Determinants
A.6.2 Elliptical Potential Lemma
@@ A.6.3 UCB Optimisation
$=\hat{z}_{t}-1^{\top}\psi(s_{t},z)+\max_{\left\lVert{u}\right\rVert_{V_{t-1}}\leq\beta_{t}}u^{\top}\psi(s_{t},z)$ (119)
$\leq\hat{z}_{t}-1^{\top}\psi(s_{t},z)+\max_{\left\lVert{u}\right\rVert_{V_{t-1}}\leq\beta_{t}}\left\lVert{u}\right\rVert_{V_{t-1}}\left\lVert{\psi(s_{t},z)}\right\rVert_{V_{t-1}^{-1}}$ (120)
$=\hat{z}_{t}-1^{\top}\psi(s_{t},z)+\beta_{t}\left\lVert{\psi(s_{t},z)}\right\rVert_{V_{t-1}^{-1}}$ (121)
@@ B.2 Requesting Reward Labels Explicitly
This approach can be directly applied to OpTI-BFM by introducing a hyper-parameter $\kappa$ and only requesting $r_{t}$ if $\Delta(\phi_{t})\geq\kappa$. We evaluate labeling efficiency for this variant: for each labeling budget $n$ (# Samples), we measure average rewards at the time step $t$ in which the budget is exhausted, i.e. $\tfrac{1}{t}\sum_{i=1}^{t}r_{t}$. This is presented in Fig.˜7 (see also Fig.˜19 and Fig.˜20 for per-task results). We observe that $\kappa$ trades-off environment interaction and labeling cost: for higher $\kappa$, good performance is achieved with fewer labels, but more episodes are required to request the same number of labels. Given a specific cost for labels, a slightly more involved thresholding scheme (Tucker et al., 2023) could potentially be used to decide the threshold apriori. Surprisingly, this variant of OpTI-BFM can maintain the performance of the main algorithm, especially in easier tasks, while reducing the amount of labeled data by one order of magnitude.
@@
Figure 7: Average reward in Cheetah at different numbers of requested reward labels (# Samples) of our main methods for different information thresholds $\kappa$. We stop interaction after 30k environment steps. $\kappa$ trades-off interaction cost with labeling cost. More than one order of magnitude less reward labels can still result in the same performance in easier tasks! We report per method results of all environments in Fig.˜20 and Fig.˜19
B.3 First-order Optimization of the UCB Objective
This section introduces a variant of OpTI-BFM which uses gradient ascent to optimize the UCB objective:
$\operatorname*{arg\,max}_{z\in\mathcal{Z}}\,\psi(s_{t},z)^{\top}\hat{z}_{t-1}+\beta_{t}\left\lVert{\psi(s_{t},z)}\right\rVert_{V_{t-1}^{-1}},$ (124)
as opposed to the random shooting approach we describe in Section˜4. We validate this variant empirically on walker tasks, relying on the Adam (Kingma & Ba, 2017) optimizer with a learning rate of $0.1$ to perform $2$, 44, 88, or $16$ gradient steps in each step, and maintaining the optimizer state throughout. At each step, we utilize $\hat{z}_{t-1}$ to warm-start the optimization procedure. We report overall return curves in Fig.˜8, and an ablation of compute costs in Fig.˜9. We observe that, given a sufficient number of gradient steps, this variant approaches the main instantiation of our method in performance. However, compute requirements for gradient ascent scale linearly with the number of gradient steps, and surpass those of random shooting. Considering that state embeddings are, for most BFMs, relatively low-dimensional (e.g. $d=50$), we conclude that random shooting approach should be preferred, as it is also less prone to fall into local optima. For very high-dimensional task spaces, first-order optimization remains however a promising option.
Figure 8: Relative performance after different # of environment interactions of OpTI-BFM and OpTI-BFM +grad; the latter is a variant that optimizes the USB objective through 88 gradient steps. Performances are similar; we report per-task results in Fig.˜18 in Appendix˜C.
(a) Absolute performance for different # of samples or gradient steps.
(b) Inference speed in Hz for different # of samples or gradient steps.
Figure 9: Performance and inference speed of OpTI-BFM, and a gradient-based variant (OpTI-BFM +grad). The performance gap is moderate, but the computational cost of the gradient-based variant is generally higher.
B.4 OpTI-BFM for Loco-navigation
To evaluate effectiveness beyond the locomotion experiments presented in the main text, we evaluate OpTI-BFM in antmaze-medium-navigate-v0, from the OGBench (Park et al., 2024a) benchmark. By default, this environment defines an indicator function reward that is 1 if the piloted quadruped reaches a certain goal position, at which point the episode terminates immediately. To make this environment consistent with our theory, we change the task to “holding” the goal position: the episode does not terminate until 1000 steps are accumulated. Furthermore, we slightly increase the radius of the sparse reward to cover the maze cell containing the goal, as seen in Fig.˜10(b) in green, as reaching the goal exactly is otherwise challenging during online learning. As is common in this benchmark (Park et al., 2024a), we introduce a behavior cloning coefficient of $0.01$ to the actor loss in pre-training, as the pre-training data contains near-expert navigation behavior as opposed to the high-coverage exploration data in ExORL. Results shown in Fig.˜10 are consistent with our main evaluation: OpTI-BFM can reach near-Oracle performance within a handful of episodes.
(a) Success Rate over 30 episodes.
(b) The first 9 trajectories (left-to-right) of Oracle and OpTI-BFM in Task 2. The initial state is in the top left (red) and the goal is to reach and stay in the cell in the bottom right (green), which assigns a reward of 1 at each step. OpTI-BFM explores the whole maze before finally exploiting the reward it has seen in the green cell. We attribute any perceived exploration behavior of the Oracle to the high behavior cloning coefficient that is necessary to train policies that perform well in this setting.
Figure 10: OpTI-BFM can infer a goal-reaching task in the OGBench (Park et al., 2024a) loconavigation environent antmaze-medium-navigate-v0.
B.5 OpTI-BFM with Inaccurate Successor Features
To investigate how robust OpTI-BFM is to violations of assuption A1, we evaluate its performance when introducing a systematic mismatch between the discounted sum of $\phi$ and $\psi$. Concretely, we let OpTI-BFM use
$\psi^{\prime}(s_{t},z)=\psi(s_{t},z)+\alpha\cdot\text{MLP}(z;\theta)$ (125)
where MLP is a two layer network with ReLU activation and L2-normalized outputs, and $\theta$ are randomly sampled weights, which we resample at the start of task inference. This additional MLP introduces a systematic, $z$-dependent bias, whose magnitude may be controlled through the hyperparameter $\alpha$. To avoid confounding factors as much as possible, we adopt an optimization procedure that samples $z$ from the whole $\mathcal{Z}$ to optimize the UCB objective, increasing the number of samples to $n=512$. As expected, Fig.˜11 shows that performance deteriorates as we increase $\alpha$; very large values of $\alpha$ ($\approx 1000$) are necessary to render OpTI-BFM completely uninformative, and approach performance of the random baseline. As the norm of $\psi$ is generally bounded by $\frac{1}{1-\gamma}\sqrt{d}\approx 350$, these constitute significant perturbations.
Figure 11: Performance of OpTI-BFM over 10 episodes in the Cheetah environment with a mismatch between $\psi$ and $\phi$ (see Eq.˜125) of magnitude $\alpha$. Performance deteriorates as $\alpha$ increases.
B.6 OpTI-BFM with Non-linear Rewards
This section investigates how a violation of assumption A2, i.e. linearity of rewards in the features $\phi$, impacts performance of OpTI-BFM. In order to isolate this effect as much as possible from other factors (e.g., how hard the task specified by a reward function is), we consider a family of reward functions with increasingly larger orthogonal components to the feature space. Starting from an existing reward function $r(s)$, we project it to the feature space, obtaining $z_{r}$, and then extract its orthogonal component $e(s)=r(s)-\phi(s)^{\top}z_{r}$ by training a small network. We can then control linearity of rewards and investigate performance under the reward function
$r_{\alpha}(s)=z_{r}^{\top}\phi(s)+\alpha\cdot e(s).$ (126)
Note that $r_{0}(s)=z_{r}^{\top}\phi(s)$ which has 0 projection error, and $r_{1}(s)$ recovers $r(s)$. While this technique can approximately disentangle reward components, for $\alpha>1$ the reward function might increase in scale: episodic performance might thus actually grow with projection error. Furthermore, the function approximation of $e$ might be inaccurate, meaning that $\alpha$ does not allow directly tuning the projection error. We can nonetheless compare the return of OpTI-BFM (upon convergence, i.e. at its 10th episode) with that of the Oracle, over a range of mean absolute projection errors, as Fig.˜12. While the results are increasingly noisy, as the error increases far beyond ranges encountered for existing rewards, OpTI-BFM seems to suffer slightly more in specific tasks—generally achieving worse performance compared to the Oracle for similarly misaligned reward functions. This may be explained by the fact that OpTI-BFM relies on linearity for both data collection and task inference, while the Oracle relies on independently collected data.
Figure 12: Return of the 10th episode of Oracle and OpTI-BFM in the Cheetah environment when artificially increasing or decreasing the projection error of the reward function onto $\phi$ through a learned network.
B.7 Can OpTI-BFM learn from noisy rewards?
In this section we evaluate the performance of OpTI-BFM under noisy reward feedback. To this end, we add zero-mean Gaussian noise to the observed rewards with different standard deviations $\sigma$. As predicted by Section˜A.4, we see in Fig.˜13 that convergence is slower for higher standard deviations. Adjusting the hyper-parameter $\beta=10$ for $\sigma=10$ did improve performance slightly, but with a noise level that is 10x higher than the reward of the environment $r_{t}\in[-1,1]$ convergence remains slow.
Figure 13: Performance of OpTI-BFM over 10 episodes in the Cheetah environment when adding Gaussian noise with different standard deviations to the rewards it observes. Convergence slows down as the noise is increased.
@@ Appendix C Full Experimental Results
(a) DMC Walker
(b) DMC Cheetah
(c) DMC Quadruped
Figure 18: Return of OpTI-BFM and OpTI-BFM +grad. that use random-shooting and gradient ascent respectively to optimize the UCB objective. The two variants perform very similarly.
@@ Appendix E Experiment Protocol
We evaluate our methods on a common zero-shot RL benchmark (Touati et al., 2023; Agarwal et al., 2025b). The benchmark consists of the Walker, Cheetah, and Quadruped environments from DMC (Tassa et al., 2018) with four tasks (reward functions) each. Note that each task has randomized initial states, so when evaluating single episode performance, e.g. Fig.˜15, we report the mean over 20 episodes. And when evaluating task inference performance over multiple episode, e.g., Figs.˜2 and 16, we report the mean over 10 trials. All error-bars and shaded regions denote min-max-intervals over three training seeds around mean performance.
@@ F.1 OpTI-BFM
We found both OpTI-BFM and OpTI-BFM-TS to be very robust to the hyper-parameters we tested. For each method we consider a single hyper-parameter with three values each: For UCB, we test a fixed $\beta_{t}=\beta\in\{1.0,0.1,0.001\}$. For TS, we test $\sigma\in\{0.1,0.001,0.0001\}$. This means all other parameters where held constant. Specifically, $\rho=1$ and $\lambda=1$ if not reported otherwise (Fig.˜5). The number of samples for the UCB optimization was $n=128$ throughout if not specified otherwise.
@@ F.2 LoLA
  • horizon length and task embedding update rate $\{50,100,250\}$ as in (Sikchi et al., 2025);
  • learning rate $\{0.1,0.05\}$ as in (Sikchi et al., 2025);